IhelpUuHELPme:

2 × 3 exponent 2 + 2 ÷ 2 - (5 + 3)

IhelpUuHELPme:

by the way it is 3 exponent 2

darkknight:

Why is this in photography?

darkknight:

You want to perform using order of operations, then u wanna solve

snowflake0531:

\[2(3^2)+2 \div 2 -(5+3) = 2(9) + 1 -8 = 18+1-8 = 11\]

darkknight:

?

snowflake0531:

:) 😔🤧

IhelpUuHELPme:

flakes ._.

snowflake0531:

hmmmm? XDD

IhelpUuHELPme:

5- months- ago-

snowflake0531:

i answered another 😔🤧😔🤧😔🤧